# MGDepth: Motion-Guided Cost Volume For Self-Supervised Monocular Depth In Dynamic Scenarios

### v1

@@ 3.5 Attention-Based Depth Network
However, instead of exclusively utilizing the feature map from the last stage of each branch for disparity map prediction, we leverage the efficient channel attention mechanism. This mechanism integrates feature maps from the current branch’s various stages, feature maps from deeper branches, and disparities from deeper branches. This fusion process can be expressed as $\mathcal{A}([f^{b}_{s}]_{s=b}^{s=4},[f^{b+1}_{s}]_{s=b+1}^{s=4},D^{b+1})$, as in Fig. 6 for the $b^{th}$ branch, to collectively inform the disparity estimation process.
3.6 Loss Function
@@
To summarize, there are three loss functions used in updating MGDepth’s weights. The final loss function $\mathcal{L}$ could be written as:
@@
where $\mathcal{M}_{o}$ is the auto mask introduced in godard2017unsupervised . $\mathcal{L}_{s}$ is the smooth loss, which could be written as:
@@
where $d_{t}^{:}*=d_{t}/\overline{d_{t}}$ and $\overline{d_{t}}$ is the average depth of target frame wang2018learning . And $\mathcal{L}_{c}$ is the consistency loss introduced by watson2021temporal , to preserve the consistency between prior monocular depth and final multi-frame depth, which could be written as:
@@
3.7 Network Structure
3.7.1 Depth Prior Network
The architecture of the Depth Prior Network (DPN) encompasses a dual-component system consisting of a depth encoder and a depth decoder. The depth encoder is established employing the HRNet16 model wang2020deep , which has been pre-trained on the ImageNet dataset deng2009imagenet and functions as the foundational backbone. This encoder operates across four distinct branches, each corresponding to a distinct pyramidal scale. The feature representations within each of these scales are subsequently channeled to the depth decoder through a sequence of skip connections, notably influenced by the UNet framework ronneberger2015u .
The depth decoder is characterized by an attention-based architecture as originally introduced in our primary manuscript. This component is designed to facilitate the fusion of encoded features derived from the depth encoder.
3.7.2 Pose Network
PoseNet exclusively comprises an encoder architecture utilizing a ResNet18 model he2016deep that has been pre-trained on the ImageNet dataset deng2009imagenet . Nonetheless, its distinctive attribute lies in its capacity to generate a 6-degree-of-freedom camera ego-motion vector, including a 3d translational vector and a 3d rotational vector.
3.7.3 Depth Network
In contrast to the Depth Prior Network, the depth network has a very similar layout. The key distinction is that, rather than analyzing just one frame at a time, it often processes two consecutive frames together. Once the first stage’s feature extraction is complete, these paired feature maps are employed to create a ”cost volume” as outlined in the main text. This cost volume is then blended with the feature representation of the target frame, resulting in a comprehensive ”final volume.” This final volume plays a crucial role in the subsequent process of building the depth map.
3.7.4 Optical Flow Network
The low-level computer vision community has favored the task of estimating the pixel-wise correspondence (Optic Flow) between two frames due to its applicability across various scenarios, without the need for specialized prior knowledge. Our paper involves adapting the pre-trained Gmflow xu2022gmflow on synthetic datasets, including the Flying Thing, Monkaa, and Driving datasets from the Scene Flow Dataset MIFDB16 and the Sintel dataset butler2012naturalistic . Notably, Gmflow used in MGDepth was not trained on any dataset used in our depth and ego-motion estimation experiments. The rationale behind adopting Gmflow lies in its lightweight nature and rapid inference capabilities, attributed to the inclusion of a dilatation correlation layer. This selection was made based on the consideration that integrating Gmflow into the existing depth estimation framework would not impose substantial computational overhead.
Figure 7: Qualitative results on KITTI Dataset. The initial row displays the RGB images of target frames where depth information has been estimated. The subsequent rows, specifically the second, third, and fourth, showcase depth maps generated by Monodepth2 trained with stereo techniques, ManyDepth, and MGDepth, respectively.
3.8 Discussion And Thinking
MGDepth distinguishes itself from previous methods, such as feng2022disentangling , that utilize category information for segmentation and object tracking. Unlike those methods, MGDepth operates independently of any prior dataset knowledge. The optical flow module, pretrained solely on synthetic testing data, significantly enhances its ability to adapt to various scenarios, eliminating the need for extra labeling efforts.
Furthermore, unlike typical motion segmentation modules relying on optical flow yang2021learning that demand substantial computational resources, MGDepth takes a different approach. We directly leverage the optical flow model and integrate necessary information, resulting in reduced computational demands.
4 Experiment
The primary experiments were conducted on the KITTI-2015 geiger2012we and Cityscapes silberman2012indoor datasets. (1) KITTI-2015: We use the Eigen split eigen2015predicting for both training and testing our model. To maintain consistency with prior studies zhou2017unsupervised ; zhou2021r , we utilized the same data selection approach and excluded all static images from the training set. This resulted in 39,810 training images and 4,424 test images.
To ensure uniformity, we employed the camera matrix configuration described in godard2019digging , which applies a single intrinsic matrix to all images and sets the length focal to the average value.
(2) Cityscapes: To ensure consistency with previous investigations watson2021temporal ; yang2018lego , we selected 69,731 images from the monocular sequences from Cityscapes cordts2016cityscapes as our training dataset, which were preprocessed into triples using scripts from zhou2017unsupervised . Our approach does not take into account stereo pairs or semantics. Performance assessment was based on the 1,525 test images, and the provided SGM hirschmuller2007stereo disparity maps were used. Similar to KITTI, we restricted depths to a max of 80m.
@@ 4 Experiment
KITTI 2015
Stereo
EPC++luo2019every 1024 x 320 0.128 0.935 5.011 0.209 0.831 0.945 0.979
FeatDepthshu2020feature 1024 x 320 0.099 0.697 4.427 0.184 0.889 0.963 0.982
R-MSFM6zhou2021r 1024 x 320 0.108 0.753 4.469 0.185 0.888 0.963 0.982
DFRzhan2018unsupervised 608 x 160 0.135 1.132 5.585 0.229 0.820 0.933 0.971
Monocular
Packnet-SFM guizilini20203d 1280 x 384 0.107 0.802 4.538 0.186 0.889 0.962 0.981
Monodepth2 godard2019digging 1024 x 320 0.115 0.882 4.701 0.190 0.879 0.961 0.982
Zhao zhao2020towards $\bullet$ 1024 x 320 0.113 0.704 4.581 0.184 0.871 0.961 0.984
Chen chen2022compacted $\bullet$ $\bullet$ 1024 x 320 0.112 0.859 4.798 0.193 0.879 0.960 0.981
FeatDepth shu2020feature 1024 x 320 0.104 0.729 4.481 0.179 0.893 0.965 0.984
Wang wang2020self $\bullet$ 1024 x 320 0.106 0.773 4.491 0.185 0.890 0.962 0.982
DevNet zhou2022devnet 1024 x 320 0.103 0.713 4.459 0.177 0.890 0.965 0.982
Ranjan ranjan2019competitive $\bullet$ 832 x 256 0.148 1.149 5.464 0.226 0.815 0.935 0.973
EPC++ luo2019every 832 x 256 0.141 1.029 5.350 0.216 0.816 0.941 0.976
Guizilini guizilini2020semantically $\bullet$ 640 x 192 0.102 0.698 4.381 0.178 0.896 0.964 0.984
Johnston johnston2020self 640 x 192 0.106 0.861 4.699 0.185 0.889 0.962 0.982
Monodepth2 godard2019digging 640 x 192 0.115 0.903 4.863 0.193 0.877 0.959 0.981
Packnet-SFM guizilini20203d 640 x 192 0.111 0.785 4.601 0.189 0.878 0.960 0.982
Patil patil2020don $\bullet$ $\bullet$ 640 x 192 0.111 0.821 4.650 0.187 0.883 0.961 0.982
Wang wang2020self $\bullet$ 640 x 192 0.106 0.799 4.662 0.187 0.889 0.961 0.982
Li li2023learning $\bullet$ $\bullet$ 640 x 192 0.102 0.703 4.348 0.175 0.895 0.966 0.984
ManyDepthwatson2021temporal $\bullet$ 640 x 192 0.098 0.770 4.459 0.176 0.900 0.965 0.983
Lee lee2023multi $\bullet$ 640 x 192 0.096 0.644 4.230 0.172 0.903 0.968 0.985
DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
MGDepth $\bullet$ $\bullet$ 640 x 192 0.091 0.650 4.263 0.171 0.908 0.967 0.984
Cityscapes
Monocular
Lee lee2021attentive 832 x 256 0.116 1.213 6.695 0.186 0.852 0.951 0.982
InstaDM lee2021learning $\bullet$ 832 x 256 0.111 1.158 6.437 0.182 0.868 0.961 0.983
Pilzer pilzer2018unsupervised 512 x 256 0.240 4.264 8.049 0.334 0.710 0.871 0.937
Monodepth2 godard2019digging 416 x 128 0.129 1.569 6.876 0.187 0.849 0.957 0.983
Videos in the Wild gordon2019depth 416 x 128 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li li2021unsupervised 416 x 128 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Struct2Depth casser2019depth $\bullet$ 416 x 128 0.151 2.492 7.024 0.202 0.826 0.937 0.972
ManyDepthwatson2021temporal $\bullet$ 416 x 128 0.114 1.193 6.223 0.170 0.875 0.967 0.989
DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 416 x 128 0.103 1.000 5.867 0.157 0.895 0.974 0.991
MGDepth $\bullet$ $\bullet$ 416 x 128 0.097 0.792 5.827 0.154 0.903 0.975 0.993
@@
Table 1: Self-supervised depth estimation results on the KITTI 2015 and Cityscapes. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The upper portion shows results on the KITTI 2015 dataset and the lower portion shows results on the Cityscapes dataset. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow.
(3) KITTI Odometry: For the purpose of odometry evaluation, we subject our system to the official KITTI odometry split, encompassing 11 driving sequences. These sequences are accompanied by ground truth odometry acquired through IMU and GPS readings. In line with prior research endeavors zhou2017unsupervised ; zhou2021r , our model’s training is conducted on sequences 00-08, while sequences 09-10 are reserved for testing.
The primary focus of this section is to evaluate the ability of MGDepth and other state-of-the-art methods to predict depth information and estimate ego-motion. The depth prediction results are evaluated using widely used metrics godard2019digging ; eigen2014depth .
The initial evaluation involves assessing MGDepth’s depth estimation performance on the KITTI-2015 and Cityscapes. Next, we evaluate its ability to estimate depth for dynamic objects on Cityscapes, utilizing the ground truth motion segmentation map. Following this, we conduct ablation studies to investigate the effectiveness of each module proposed in our paper. Finally, we present the results of MGDepth’s ego-motion estimation performance on the KITTI Odometry dataset.
4.1 Monocular Depth Estimation
@@
KITTI-2015: In this experiment, the ResNet18 he2016deep architecture was utilized as the backbone for the pose network. The HRNet16 was used as the backbone for coarse depth network and attention-based depth network. During the training phase, our MGDepth receives input images with a resolution of $192\times 640$. For testing, we set the minimum and maximum depth values to 0.1m and 80m, respectively, following the recommendations of godard2019digging ; jin2001real . In our experiment, the scale factor is computed using the median value of the ground-truth image, employing the same approach as in zhou2017unsupervised . The upper section of Tab. 1 presents a comparison between the performance of MGDepth and state-of-the-art algorithms. Overall, MGDepth exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used. In particular, MGDepth demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth. MGDepth outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel. In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, MGDepth achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information. Quantitative results are shown in Fig. 7.
@@
Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects. During the training phase on the Cityscapes dataset, the input resolution of MGDepth is set to $128\times 416$. The lower section of Tab. 1 indicates a substantial performance advantage of MGDepth over its competing methods. Remarkably, MGDepth exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric. Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig. 8 facilitates a qualitative analysis of depth predictions. These observations indicate that MGDepth offers enhanced accuracy for depth estimation of both dynamic foreground and static background. Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for MGDepth. Quantitative results are shown in Fig. 8.
Type Method M WxH Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
S InstaDM lee2021learning 832 x 256 0.139 1.698 5.760 0.181 0.859 0.959 0.982
w/o S Monodepth2 godard2019digging 416 x 128 0.159 1.937 6.363 0.201 0.816 0.950 0.981
ManyDepth watson2021temporal $\bullet$ 416 x 128 0.169 2.175 6.634 0.218 0.789 0.921 0.969
MGDepth $\bullet$ 416 x 128 0.123 1.260 4.519 0.144 0.869 0.971 0.989
@@
Table 2: Results of Dynamic Foreground on Cityscapes. We evaluate depth prediction errors for dynamic objects such as vehicles, people, and bikes on the Cityscapes dataset silberman2012indoor . The best results are highlighted in bold.
Figure 8: Qualitative results on Cityscapes (Error Map). The first row exhibits the RGB image of target frames for which depth information is estimated. The second and third rows illustrate error maps depicting the absolute relative error concerning the ground truth, where the degree of accuracy is represented by white (good) to red (poor) color gradations. The mask proposed in watson2021temporal is employed during drawing.
4.2 Depth For Dynamic Foreground
Tab. 2 showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in lee2021learning , which corresponds to the qualitative results presented in Fig. 8. Based on these results, it can be inferred that self-supervised monocular depth estimation methods from previous studies struggle to accurately estimate depth in dynamic scenes. This is likely due to their reliance on photometric loss, which assumes that the majority of foreground objects are static. In dynamic scenes, the photometric loss assumption is often violated, leading to difficulty in providing accurate depth estimates. However, MGDepth presents an effective solution to this issue by handling dynamic objects. Moreover, the motion-guided ego-motion loss function used in MGDepth enables accurate pose estimation, which aids in constructing cost volumes similar to stereo depth estimation. As a result, MGDepth achieves superior performance.
Figure 9: Trajectory visualization on Seq. 09. and Seq. 10. The ground truth trajectories are represented by red lines. Trajectories generated by the FeatDepth shu2020feature results are indicated by green lines, while our method’s trajectories are depicted in blue lines.
@@ 4.2 Depth For Dynamic Foreground
DFR zhan2018unsupervised 11.93 3.91 12.45 3.46
MonoDepth2 godard2019digging 10.85 2.86 11.60 5.72
NeuralBundler li2019pose 8.10 2.81 12.90 3.71
SC-SfMlearner zhou2017unsupervised 8.24 2.19 10.70 4.58
FeatDepth shu2020feature 8.73$\dagger$ 2.11 10.65$\dagger$ 4.91
ManyDepthwatson2021temporal 8.08 1.97 9.86 3.42
MGDepth 7.01 1.76 7.29 2.65
Table 3: Visual odometry results on Seq. 9 and Seq. 10 of the KITTI odometry dataset. The average drift in root mean square error for both translation (Tr) and rotation (R) is reported. $\dagger$ denotes the method that outperforms the results reported in the original paper.
4.3 Odometry Estimation
To evaluate odometry estimation results, we follow the split used in zhou2017unsupervised ; zhan2018unsupervised . Specifically, we trained our model using Seq. 00-08 from the KITTI odometry dataset and test methods on Seq. 09-10. Tab. 3 shows the results, where the translational and rotational motion were evaluated by using the root mean square error (RMSE). MGDepth, which uses less computational resources, outperforms other learning-based techniques that use more computational resources, such as FeatDepth. The results indicate that MGDepth achieves a significant reduction in RMSE by approximately $24.1\%$ for translational motion and $22.5\%$ for rotational motion, as compared to ManyDepth, on Seq. 10. In support of this result, we present a visual depiction of the trajectory of both methods in Fig. 9, which shows that MGDepth’s trajectory exhibits considerably less drift than that of FeatDepth. The superior performance of MGDepth in precise pose estimation can be attributed to its use of both depth-based photometric loss and flow-based ego-motion loss, as well as the implementation of a motion-guided mask that effectively filters out dynamic foreground outliers. This combination of techniques enables MGDepth to achieve more reliable pose estimates compared to others.
@@ 4.3 Odometry Estimation
Final Prior Depth 0.108 0.753 4.563
@@
MGDepth 0.091 0.650 4.263
Table 4: Ablation study results of depth accuracy on the KITTI 2015. The proposed MGDepth model was evaluated by modifying its components.
4.4 Abaltion Study
We conduct ablation studies on the Cityscapes dataset to demonstrate the effectiveness of our proposed modules and validate our design choices. The results are presented in Tab. 4, where the experiments were conducted under various settings, including Prior Depth Net of MGDepth, MGDepth without Motion-Guided Volume, MGDepth without Attention for Final Volume, and MGDepth with Normal HRNet. Tab. 4 indicates that the addition of Attention-Based HRNet has a significant impact on the performance of MGDepth. And the introduction of Motion-Guided Volume and the Attention for Final Volume could further improve the performance of MGDepth.
5 Conclusion
We present MGDepth, a self-supervised multi-frame monocular depth prediction model leveraging optical flow-depth geometry for enhanced depth estimation. Our approach involves novel components such as a static reference frame, motion-guided cost volume, and an attention-based high-resolution neural network. MGDepth achieves accurate depth estimation for dynamic foreground and static background on KITTI-2015 and Cityscapes datasets. Notably, our results are reproducible using a single NVIDIA A10 GPU. MGDepth seamlessly integrates with existing self-supervised monocular depth estimation algorithms, serving as a refined coarse depth network. Moreover, our proposed modules augment accuracy by incorporating dynamic information.
6 Acknowledgements

### v2

@@ III-E Attention-Based Depth Network
The High-Resolution Network (HRNet) [17] is well-regarded for its ability to preserve a high level of detail in input images. The HRNet is composed of multiple branches, denoted as $B$, each generating $S=b,...,4$ features with resolutions of $(\frac{H}{2^{b-1}},\frac{W}{2^{b-1}})$. However, instead of exclusively utilizing the feature map from the last stage of each branch for disparity map prediction, we leverage the channel attention mechanism. This mechanism integrates feature maps from the current branch’s various stages and feature maps from deeper branches. This fusion process can be expressed as $\mathcal{A}([f^{b}_{s}]_{s=b}^{s=4},[f^{b+1}_{s}]_{s=b+1}^{s=4})$.
III-F Loss Function
@@
To summarize, there are three loss functions used in updating Manydepth2’s weights. The final loss function $\mathcal{L}$ could be written as:
@@
where $\mathcal{M}_{o}$ is the auto mask introduced in [18]. $\mathcal{L}_{s}$ is the smooth loss, which could be written as:
@@
where $d_{t}^{:}*=d_{t}/\overline{d_{t}}$ and $\overline{d_{t}}$ is the average depth of target frame [19]. And $\mathcal{L}_{c}$ is the consistency loss introduced by [7], to preserve the consistency between prior monocular depth and final multi-frame depth, which could be written as:
@@
IV Experiments
@@
KITTI 2015 Stereo EPC++[20] 1024 x 320 0.128 0.935 5.011 0.209 0.831 0.945 0.979
FeatDepth[21] 1024 x 320 0.099 0.697 4.427 0.184 0.889 0.963 0.982
DFR[22] 608 x 160 0.135 1.132 5.585 0.229 0.820 0.933 0.971
Monocular Packnet-SFM [23] 1280 x 384 0.107 0.802 4.538 0.186 0.889 0.962 0.981
Monodepth2 [24] 1024 x 320 0.115 0.882 4.701 0.190 0.879 0.961 0.982
FeatDepth [21] 1024 x 320 0.104 0.729 4.481 0.179 0.893 0.965 0.984
DevNet [9] 1024 x 320 0.103 0.713 4.459 0.177 0.890 0.965 0.982
Ranjan [25] $\bullet$ 832 x 256 0.148 1.149 5.464 0.226 0.815 0.935 0.973
EPC++ [20] 832 x 256 0.141 1.029 5.350 0.216 0.816 0.941 0.976
Guizilini [26] $\bullet$ 640 x 192 0.102 0.698 4.381 0.178 0.896 0.964 0.984
Monodepth2 [24] 640 x 192 0.115 0.903 4.863 0.193 0.877 0.959 0.981
Packnet-SFM [23] 640 x 192 0.111 0.785 4.601 0.189 0.878 0.960 0.982
Patil [11] $\bullet$ $\bullet$ 640 x 192 0.111 0.821 4.650 0.187 0.883 0.961 0.982
Li [27] $\bullet$ $\bullet$ 640 x 192 0.102 0.703 4.348 0.175 0.895 0.966 0.984
ManyDepth[7] $\bullet$ 640 x 192 0.098 0.770 4.459 0.176 0.900 0.965 0.983
Lee[28] $\bullet$ 640 x 192 0.096 0.644 4.230 0.172 0.903 0.968 0.985
AQUANet[29] $\bullet$ 640 x 192 0.105 0.621 4.227 0.179 0.889 0.964 0.984
DCPI-Depth[30] $\bullet$ $\bullet$ 640 x 192 0.095 0.662 4.274 0.170 0.902 0.967 0.985
DynamicDepth [15] $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
Manydepth2 $\bullet$ $\bullet$ 640 x 192 0.091 0.649 4.232 0.170 0.909 0.968 0.984
@@
TABLE I: Self-supervised depth estimation results on the KITTI 2015. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow.
Cityscapes Monocular Lee [31] 832 x 256 0.116 1.213 6.695 0.186 0.852 0.951 0.982
InstaDM [16] $\bullet$ 832 x 256 0.111 1.158 6.437 0.182 0.868 0.961 0.983
Pilzer [32] 512 x 256 0.240 4.264 8.049 0.334 0.710 0.871 0.937
Monodepth2 [24] 416 x 128 0.129 1.569 6.876 0.187 0.849 0.957 0.983
Videos in the Wild [33] 416 x 128 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li [34] 416 x 128 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Struct2Depth [10] $\bullet$ 416 x 128 0.151 2.492 7.024 0.202 0.826 0.937 0.972
ManyDepth[7] $\bullet$ 416 x 128 0.114 1.193 6.223 0.170 0.875 0.967 0.989
DynamicDepth [15] $\bullet$ $\bullet$ 416 x 128 0.103 1.000 5.867 0.157 0.895 0.974 0.991
Manydepth2 $\bullet$ $\bullet$ 416 x 128 0.097 0.792 5.827 0.154 0.903 0.975 0.993
TABLE II: Self-supervised depth estimation results on the Cityscapes. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow.
The main tests and evaluations for our study were carried out using two primary datasets: KITTI-2015 [35] and Cityscapes [36]. In our experiments, the ResNet18 [37] architecture was utilized for the pose network. The HRNet16 was used as the backbone for the attention-based depth network. The pre-trained Gmflow [38] on the Sintel dataset [39] are used as optic flow network. For testing, we set the minimum and maximum depth values to 0.1m and 80m, as in [24, 40]. In our experiment, the scale factor is computed using the median value of the ground-truth image as in [8].
IV-A Monocular Depth Estimation
@@
KITTI-2015: The Tab. I presents a comparison between the performance of Manydepth2 and state-of-the-art algorithms. Overall, Manydepth2 exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used. In particular, Manydepth2 demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth. Manydepth2 outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel. In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, Manydepth2 achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information. Quantitative results are shown in Fig. 4.
@@
Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects. During the training phase on the Cityscapes dataset, the input resolution of Manydepth2 is set to $128\times 416$. The lower section of Tab. II indicates a substantial performance advantage of Manydepth2 over its competing methods. Remarkably, Manydepth2 exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric. Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig. 1 facilitates a qualitative analysis of depth predictions. These observations indicate that Manydepth2 offers enhanced accuracy for depth estimation of both dynamic foreground and static background. Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for Manydepth2. Quantitative results are shown in Fig. 1. Tab. III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results presented in Fig. 1.
Type Method M WxH Abs Rel Sq Rel RMSE RMSE log
S InstaDM [16] 832 x 256 0.139 1.698 5.760 0.181
w/o S Monodepth2 [24] 416 x 128 0.159 1.937 6.363 0.201
ManyDepth [7] $\bullet$ 416 x 128 0.169 2.175 6.634 0.218
Manydepth2 $\bullet$ 416 x 128 0.123 1.260 4.519 0.144
@@
TABLE III: Results of Dynamic Foreground on Cityscapes. We evaluate depth prediction errors for dynamic objects such as vehicles, people, and bikes on the Cityscapes dataset [36]. The best results are highlighted in bold.
@@
DFR [22] 11.93 3.91 12.45 3.46
NeuralBundler [41] 8.10 2.81 12.90 3.71
ManyDepth[7] 8.08 1.97 9.86 3.42
Manydepth2 7.01 1.76 7.29 2.65
TABLE IV: Visual odometry results on Seq. 9 and Seq. 10 of the KITTI odometry dataset. The average drift in root mean square error for both translation (Tr) and rotation (R) is reported.
@@
Final Prior Depth 0.105 0.768 4.536
@@
Manydepth2 0.091 0.650 4.263
TABLE V: Ablation study results of depth accuracy on the KITTI 2015. The proposed Manydepth2 model was evaluated by modifying its components.
IV-B Odometry Estimation
To evaluate odometry estimation results, we follow the split used in [8, 22]. Specifically, we trained our model using Seq. 00-08 from the KITTI odometry dataset and test methods on Seq. 09-10. Tab. IV shows the results, where the translational and rotational motion were evaluated by using the root mean square error (RMSE). Manydepth2, which uses less computational resources, outperforms other learning-based techniques that use more computational resources, such as FeatDepth. The results indicate that Manydepth2 achieves a significant reduction in RMSE by approximately $24.1\%$ for translational motion and $22.5\%$ for rotational motion, as compared to ManyDepth, on Seq. 10. In support of this result, we present a visual depiction of the trajectory of both methods in Fig. 5, which shows that Manydepth2’s trajectory exhibits considerably less drift than that of FeatDepth. The superior performance of Manydepth2 in precise pose estimation can be attributed to its use of both depth-based photometric loss and flow-based ego-motion loss, as well as the implementation of a motion-guided mask that effectively filters out dynamic foreground outliers. This combination of techniques enables Manydepth2 to achieve more reliable pose estimates compared to others.
Figure 5: Trajectory visualization on Seq. 09. and Seq. 10. The ground truth trajectories are represented by red lines. Trajectories generated by the FeatDepth [21] results are indicated by green lines, while our method’s trajectories are depicted in blue lines.
IV-C Abaltion Study
In our research, we meticulously carry out ablation studies using the Cityscapes dataset. These studies are designed to dissect and showcase how effective our proposed modules are, allowing us to validate the specific design choices we’ve made. The detailed outcomes of these experiments are showcased in Table V. Within this table, we provide insights from experiments conducted under diverse conditions and configurations. Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume, without Attention for Final Volume, and with the inclusion of a Normal HRNet. A key observation from Table V is the pronounced positive impact that integrating the Attention-Based HRNet has on enhancing Manydepth2’s performance. Furthermore, our findings underscore that the incorporation of both the Motion-Guided Volume and Attention for the Final Volume amplifies the capabilities of Manydepth2 even further, leading to improved outcomes.
V Conclusion
We introduce Manydepth2, an innovative self-supervised multi-frame monocular depth prediction model that harnesses the synergy of optical flow-depth geometry. One of the standout features of Manydepth2 is its unique components, including a static reference frame, motion-guided cost volume, and an attention-enhanced high-resolution neural network. When benchmarked against the KITTI-2015 and Cityscapes datasets, Manydepth2 consistently delivers precise depth estimation, effectively distinguishing between dynamic foreground elements and static backgrounds. Remarkably, our model’s performance is achieved using just a single NVIDIA A10 GPU, emphasizing its efficiency. Furthermore, Manydepth2 not only complements existing self-supervised monocular depth estimation techniques but also elevates them by integrating dynamic insights, thereby significantly enhancing accuracy.
VI Acknowledgements